zeros of -3,1, and 4;f(2)=30

a=-3 b=1 c=4
f(x)=(x+3)(x-1)(x-4)

Answer by ikleyn(52776)

.
zeros of -3,1, and 4;f(2)=30

a=-3 b=1 c=4
f(x)=(x+3)(x-1)(x-4)
----------------------------------

f(2) = (2+3)*(2-1)*(2-4) = 5*1*(-2) = -10.

So, apply the factor (-3) to f(x):

F(x) = -3*f(x).